# ProactivePIM: Accelerating Weight-Sharing Embedding Layer with PIM for Scalable Recommendation System

### v4

@@ Abstract
The continuous growth of personalized recommendation system poses new challenges for model inference. Although weight-sharing algorithms have been proposed to reduce embedding table capacity, they increase memory access. Recent advancements in processing-in-memory (PIM) successfully enhance the recommendation system’s throughput by exploiting memory parallelism, but our analysis shows that such algorithms introduce CPU-PIM communication overhead into prior PIM systems, which compromise the PIM throughput. We propose ProactivePIM, a specialized memory architecture integrated with PIM technology tailored to accelerate the weight-sharing recommendation system. ProactivePIM integrates an SRAM cache within the PIM with an efficient prefetching scheme to leverage an unique locality of the algorithm and eliminate CPU-PIM communication overhead through a subtable mapping strategy. Our design achieves a 4.8$\times$ speedup compared to prior works.
@@ I Introduction
Deep learning-based personalized recommendation system serves as a pivotal technology in various companies [2], [16], [22], with the Deep Learning Recommendation System (DLRM) being a representative model. These systems demand substantial data center resources and account for a significant portion of the computational workload of data centers [6]. As shown in Figure 1 (a), these systems input user information (dense features) and associated items (sparse features) into the bottom-MLP and embedding layer, with the latter performing gather-and-reduce (GnR) operations. The outputs of these layers undergo feature interaction and top-MLP, which generates Click-Through-Rate (CTR).
Unfortunately, as depicted in Figure 1 (b), the memory-intensive nature of the embedding layer grows continuously, placing constraints on traditional data center infrastructures [27]. Recent studies proposed leveraging near-memory processing (NMP) and processing-in-memory (PIM) to accelerate deep learning algorithms by exploiting memory parallelism [7], [10], [11], [12], [17]. For recommendation systems, solutions like TensorDIMM [11], RecNMP [10], and TRiM [18] accelerate GnR, while SPACE [9] employs HBM-based NMP to improve power efficiency over DIMM-based designs.
@@
As shown in Figure 1 (b), the model size of DLRM has grown by 16 times from 2020 to 2024, as incorporating more items for better model quality [14], [20]. Serving inference with a multi-node server or introducing solid-state drives (SSDs) could be solutions. Still, they produce high synchronization costs, vulnerability to failures, and a negative impact on execution time [1], [28]. An algorithmic approach, the hashing trick [24], fits the model into the memory capacity budget at the cost of accuracy degradation. Weight-sharing algorithms, which transform the original embedding table into multiple small subtables, have been proposed to compensate for the accuracy degradation of hashing-trick [21], [26]. Nevertheless, these methods amplify the off-chip memory access and are still memory-intensive. Prior NMP/PIM systems for recommendation models are ineffective because the multiplication used in these algorithms requires operands from subtables to reside in the same NMP/PIM unit, unlike simpler GnR. This limitation necessitates CPU-PIM communication, causing severe overhead. Furthermore, they fail to exploit the extra data locality of these algorithms.
We propose ProactivePIM, a specialized PIM-integrated memory architecture to accelerate weight-sharing algorithm inference. ProactivePIM is built on a heterogeneous DIMM-HBM system, where a 2-level PIM system is built within the HBM: base-die PIM (bd-PIM) and bank-group PIM (bg-PIM). We analyze weight-sharing algorithms and identify two key insights: 1) Subembeddings of subtables exhibit intra-GnR locality, which can be reused multiple times within and beyond a single GnR, and 2) the smallest subtables are compact enough to fit within a bank group. To leverage these insights, we incorporate an SRAM cache in bg-PIM to prefetch subembeddings with high locality, which reduces bank access. To further reduce CPU-PIM communication overhead, we propose a subtable mapping strategy that duplicates selected subtables across bank groups. The key contributions of our work are as follows:
@@
  • We conduct a detailed analysis of weight-sharing algorithms, identifying drawbacks and intra-GnR locality.
@@ II-A Challenges in Recommendation System Inference
Industry-scale DLRM implementations reach sizes up to tens of terabytes, sometimes surpassing the maximum memory capacity budget of a single-node inference server. Given that the model quality is closely linked to the size of embedding tables [1], [27], its scale growing year by year [1], [5], [20]. However, the practical utilization of such large models is constrained by capacity limitations in the server. Furthermore, memory-intensive characteristics limit inference throughput. The sparse feature exhibits irregular and sparse memory access, making it challenging for the embedding layer to benefit from the on-chip memory, coupled with the embedding table size that surpasses the cache capacity. Therefore, the embedding layer necessitates frequent access to the off-chip memory system. As the embedding layer execution takes longer latency than the bot-MLP [9], the memory-bound nature of the embedding layer degrades the inference throughput.
@@ II-B NMP/PIM Designs for Recommendation System
Recent studies have integrated NMP/PIM units within DIMMs to accelerate the memory-intensive GnR operation of embedding layers. TensorDIMM [11] and RecNMP [10] have exploited rank-level NMP, whereas TRiM [18] leveraged bankgroup-level PIM for in-memory parallelism. These units perform partial GnR on the data allocated within its node and transfer the result to the host CPU. For embedding placement within memory, horizontal partitioning (HP) stores each embedding in a single memory node but suffers from load imbalance due to sparse, irregular access patterns. In contrast, TensorDIMM utilized vertical partitioning (VP), dividing an embedding across multiple ranks. VP is robust to the load imbalance but comes at the cost of energy efficiency as frequent row activations are incurred among multiple ranks. In pursuit of further speedup, the designs mentioned above leveraged the long-tail distribution of embeddings [9], [10] by placing a dynamic cache inside each NMP or introducing a load-balancing technique [18], [10]. SPACE enhanced throughput using HBM as a static DRAM cache for DIMMs, storing frequently accessed embeddings and partial sums [9].
@@ III-A Weight-sharing Algorithms for Embedding Table Reduction
Weight-sharing algorithms, QR-trick [21] and TT-Rec [26], improve compression ratios and mitigate the accuracy loss of the hashing trick [24]. Weight-sharing algorithms transform an embedding table into several subtables consisting of subembeddings. Therefore, they require an additional reconstruction process before the GnR. This reconstruction, termed collect-and-reconstruction (CnR) in this paper, involves element-wise multiplication or GEMV operations.
QR-trick. The QR-trick uses double hashing with quotient and remainder hash functions to generate two subtables: the Q subtable and the R subtable, as shown in Figure 2 (a). The Q subtable maps consecutive original embeddings to the same subembedding based on identical quotient values, while the R subtable maps embeddings separated by the collision value based on identical remainder values. The R subtable size is negligible compared to the Q subtable, as it equals the hash collision value. During inference, the embedding layer reconstructs the original embedding from the subtables. Each hash function processes an input index to generate Q and R subtable indices for retrieving subembeddings. A mathematical operation is then performed on these subembeddings to reconstruct the embedding vector, where this paper focuses on multiplication, which generally outperforms other operations across various models [21].
@@
TT-Rec. TT-Rec compresses embedding tables by decomposing them into tensor-train cores (TT-cores) through reshaping and decomposition, as shown in Figure 2 (b). The embedding table $T\in\mathbb{R}^{M\times N}$ is reshaped by factorizing $M$ and $N$ into $M=m1\times m2\times\cdots\times md$ and $N=n1\times n2\times\cdots\times nd$, resulting in $T\in\mathbb{R}^{(m1\times n1)\times(m2\times n2)\times\cdots\times(md\times nd)}$. Decomposition then converts the reshaped table into TT-cores, where kth TT-core shape is $C(k)=\mathbb{R}^{R_{k-1}\times(mk\times nk)\times R_{k}}$, $0<k<d$, where $R$ being the rank and $d$ the number of cores. The rank values are typically set to 16, 32, and 64 and the number of TT-cores to 3. Note that the TT-cores correspond to subtables in this work. The second subtable is three-dimensional; each entry consists of $R$ vectors. Reconstructing a full embedding is accomplished by fetching $n1$, $n2$, and $n3$ consecutive embeddings from each subtable and then performing the GEMV followed by an element-wise multiplication. TT-Rec results in a 327$\times$ reduction with the number of subtables and the rank value set to 3 and 32 [26].
@@ III-B Weight-Sharing Algorithms on Previous PIM Systems
The weight-sharing algorithm involves element-wise multiplication or GEMV operations in the embedding layer for CnR, which require vector pairs to reside in the same PIM unit for computation. However, it is impossible to co-locate all vector pairs within the same node because the subtables are distributed across multiple nodes (e.g., ranks, bank groups) for parallel processing. This requires CPU-PIM communication, as depicted in Figure 3 (a). The second CnR requires the memory controller to pass an operand in a different node, where a similar incident occurs for the TT-Rec. Furthermore, as shown in Figure 3 (b), communication overhead takes more than 30% of the total execution time of the embedding layer for all configurations. Dynamic techniques to alleviate CPU-NMP communication, such as inter-node broadcast [23], face challenges with weight-sharing algorithms. The asynchronous nature of PIM operations in the embedding layer requires each PIM to fetch distinct data, resulting in idle cycles as PIM should wait for data over the shared bus. Although interconnect methods for DIMM-NMP [29] have been proposed to address point-to-point communication, modern PIM architectures still lack support for inter-PIM communication [8]. Even without communication, weight-sharing algorithms incur increased memory access as CnR requires access to all subtables. The abovementioned problems undermine the performance of prior PIM architectures. However, our roofline analysis [25] in Figure 4 shows that both algorithms are within the memory-bound region, indicating that PIM is still a promising solution for embedding layer with weight-sharing algorithms.
@@ III-C Access Patterns and Table Characteristics
Each subtable must be accessed concurrently in weight-sharing algorithms to reconstruct an original embedding. The total number of accesses is not distributed but remains the same for each subtable. In addition, the access rate of each subembedding intensifies because the total subembeddings in a subtable are much smaller than the original embedding table. Thus, access to each subembedding is more concentrated than the original. Another observation is that subembeddings are reused not only across GnRs of the same batch but also within a single GnR. In conventional recommendation systems, a user’s interaction with an item is typically reflected once in one sparse feature, preventing multiple accesses within a single GnR. However, weight-sharing algorithms map multiple embeddings to a single subembedding, necessitating a CnR before the GnR, thus enabling the reuse of subembeddings. We term this characteristic as intra-GnR locality. As shown in Figure 2 (b), access to consecutive embeddings cause multiple access to the single subembedding within the Q subtable. The intra-GnR locality is further strengthened for smaller subtable sizes, given that access to the subembeddings gets more concentrated. Previous designs fail to capture such characteristics effectively, losing an opportunity for extra speedup.
@@ IV ProactivePIM Architecture
We propose the ProactivePIM, a PIM architecture that supports weight-sharing algorithm. ProactivePIM employs an integrated approach for speedup by in-memory cache utilization (as detailed in Section IV-B), reducing communication overhead via data mapping (as detailed in Section IV-C) while minimizing extra overhead in the PIM extension and the PIM unit.
@@ IV-A ProactivePIM Architecture
A two-level PIM system is used inside HBM to leverage in-memory parallelism. In Figure 5 (b), the first-level processing unit (bd-PIM) is on the base die, while the second-level units (bg-PIM) are inside each bank group. Each bg-PIM performs CnR followed by GnR on the CnR outputs, and passes the partial results to the bd-PIM for final reduction. Figure 5 (d) shows bg-PIM, where PIM instructions (PIM-Inst) delivered from the host are buffered and then decoded in the decoder to set control paths of the PIM unit. MMReg is placed within the decoder to indicate the prefetch address’s start and end. MAC unit performs CnR and GnR, while prefetched subembeddings are stored inside cache to utilize intra-GnR locality. The HP method is employed for the data placement, as HP’s energy efficiency is better than VP’s. Figure 5 (e) shows Bd-PIM, similar to the bg-PIM without the cache and MAC unit. ProactivePIM is also applicable to memory-intensive applications with skewed data access, such as personalized recommendation systems without weight sharing or PageRank.
@@ IV-B Hot Subembedding Prefetch
ProactivePIM leverages an integrated solution by utilizing prefetching high intra-GnR locality, removing communication overhead and minimizing extra overhead inside the memory controller. To leverage intra-GnR Locality, ”hot” subembeddings with high access frequency are cached into the SRAM of bg-PIM before the CnR/GnR execution. During the CnR/GnR, hot subembedding accesses are directed to the cache, reducing bank access. Partial sum transfer and prefetching are pipelined to minimize the prefetch overhead, by utilizing separate buses: transfer uses an external bus, while prefetching uses an internal bus, as shown in Figure 5 (d). When PIM-Inst for data transfer is issued from the decoder, bg-PIM starts to prefetch the target subtable. To ensure correct timing of PIM-Inst for cached data, storing their physical addresses in the PIM extension is necessary. ProactivePIM addresses this by prefetching an entire subtable, allowing the extension to rely solely on the subtable ID when issuing PIM-Inst for cached data access.
Target Subtable to Prefetch. The key criteria for selecting a target subtable are (i) the degree of intra-GnR locality and (ii) the size of the subtable. In QR-trick, the R subtable size is far less than the Q subtable, and subembeddings exhibit high intra-GnR locality as described in Section III-A. Therefore, the R subtable is chosen. For TT-Rec, the intra-GnR locality is similar among all the subtables, but the second subtable is three-dimensional, resulting in a larger size. Furthermore, as described in Section IV-C, the subembedding dimension of the third table is smaller than the other tables, which degrades cache efficiency. Therefore, the first subtable is the target.
Prefetch Method. Figure 6 (a) shows the original execution timeline for CnR/GnR execution of three original tables. Once CnR/GnR completes for table A, the partial sums from each bg-PIM are transferred to the bd-PIM during the transfer phase. Figure 6 (b) illustrates the naive method in which all target subtables are read and offloaded before the embedding layer execution, which we name all-table prefetch. At the cost of the initial delay, the time consumed in PIM operation is largely reduced due to the cache hit. This effect is permanent before the inference terminates. However, this method requires a large SRAM cache to store all the target subtables. When there are 40 tables, 8% area overhead in DRAM die occurs, which worsens with more tables. To mitigate this, VP could be utilized to distribute subembedding dimensions across HBM bank groups, reducing SRAM requirements fourfold. However, for subembedding dimensions smaller than 256B, common in TT-Rec, the data read per bank group falls below the DRAM burst granularity, reducing efficiency.
@@
As shown in equation 1, the transfer latency (tTransfer) is proportional to the total bank groups in a channel, read command delay between bank groups (tccd_s), batch size, and the embedding dimension (vlen). In contrast, the prefetch latency (tPrefetch) is proportional to total subembeddings within a target subtable (totsub), read command delay within bank groups (tccd_l) and subembedding dimension (subvlen). As a small batch size could cause severe load imbalance across PIM nodes, especially critical for a multi-stack HBM system, we choose a batch size 16. As described in Section VI-D, prefetch overhead is minimal, indicating that tTransfer is similar to the tPrefetch. Additionally, target subtable subembeddings are mapped consecutively in a row to utilize the row buffer hit.

### v5

@@ Abstract
The model size growth of personalized recommendation systems poses new challenges for inference. Weight-sharing algorithms have been proposed for size reduction, but they increase memory access. Recent advancements in processing-in-memory (PIM) enhanced the model throughput by exploiting memory parallelism, but such algorithms introduce massive CPU-PIM communication into prior PIM systems. We propose ProactivePIM, a PIM system for weight-sharing recommendation system acceleration. ProactivePIM integrates a cache within the PIM with a prefetching scheme to leverage a unique locality of the algorithm and eliminate communication overhead through a subtable mapping strategy. ProactivePIM achieves a 4.8$\times$ speedup compared to prior works.
@@ I Introduction
Deep learning-based personalized recommendation system serves as a pivotal technology in various companies [2], [19], [26], with the Deep Learning Recommendation System (DLRM) being a representative model. These systems account for a significant computational workload of data centers [8]. Figure 1 (a) shows the DLRM architecture, where dense features (user information) are processed through a bottom multi-layer perceptron (MLP), and sparse features (items) are handled by the embedding layer. The embedding layer performs gather-and-reduce (GnR) operations, which retrieve embedding vectors for items and perform element-wise addition. The outputs undergo feature interaction and top MLP, which generates Click-Through-Rate (CTR) as a result.
As depicted in Figure 1 (b), the memory-intensive nature of the embedding layer grows continuously, placing constraints on data center infrastructures [32]. Recent studies leveraged near-memory processing (NMP) and processing-in-memory (PIM) to accelerate AI models by exploiting memory parallelism [9], [12], [13], [14], [21]. For recommendation systems, TensorDIMM [13], RecNMP [12], TRiM [22] and SPACE [11] accelerated GnR via NMP and PIM to address memory-intensive characteristic. On the other hand, as shown in Figure 1 (b), the model size of DLRM has grown by 16 times from 2020 to 2024, as incorporating more items result in better model quality [17], [24]. Serving inference with a multi-node server or introducing solid-state drives (SSDs) could be solutions, but they produce high synchronization costs, vulnerability to failures, and a negative impact on execution time [1], [33]. An algorithmic approach, the hashing trick [28], fits the model into the memory capacity budget at the cost of accuracy degradation. Weight-sharing algorithms, which transform the original embedding table into multiple small subtables, are proposed to compensate for such drawbacks. Nevertheless, these methods amplify the off-chip memory access and are still memory-intensive. Prior NMP/PIM systems for recommendation models are ineffective because the multiplication used in these algorithms requires operands from subtables to reside in the same NMP/PIM unit, unlike simpler GnR. This limitation necessitates CPU-PIM communication, causing severe overhead. Furthermore, they are not optimized for exploiting data locality of these methods. There exist other algorithmic methods [6], [16], but the weight-sharing algorithm is superior in degree of compression.
@@
We propose ProactivePIM, a PIM-integrated memory architecture to address data locality and CPU-PIM communication of weight-sharing algorithm inference. ProactivePIM is built on a heterogeneous DIMM-HBM system, where a 2-level PIM system is built within the HBM: base-die PIM (bd-PIM) and bank-group PIM (bg-PIM). We analyze weight-sharing algorithms and identify two key insights: 1) Embeddings of subtables exhibit intra-GnR locality, which can be reused multiple times within and beyond a single GnR, and 2) the smallest subtables are compact enough to fit within a bank group. To leverage these insights, we incorporate an SRAM cache in bg-PIM to prefetch high locality embeddings before GnR execution. To reduce CPU-PIM communication overhead, we propose a subtable mapping strategy that duplicates selected subtables across bank groups. ProactivePIM can be applied for various memory-intensive workloads with skewed data access patterns and operations such as GnR or aggregation [31]. This includes applications like PageRank [7] and the original DLRM, taking advantage of its cache and 2-level PIM architecture. The key contributions of our work are as follows:
@@
  • We conduct a detailed analysis of weight-sharing algorithms, identifying its drawbacks and intra-GnR locality.
@@ II-A Challenges in Recommendation System Inference
Industry-scale DLRM implementations reach sizes up to tens of terabytes, sometimes surpassing the maximum memory capacity budget of a single-node inference server. Given that the model quality is closely linked to the size of embedding tables [1], [32], its scale is growing year by year [1], [5], [24]. However, the practical utilization of such large models is constrained by capacity limitations in the server. Furthermore, memory-intensive characteristics limit inference throughput. The sparse feature exhibits irregular and sparse memory access, making it challenging for the embedding layer to benefit from the on-chip memory, coupled with the embedding table size that surpasses the cache capacity. Therefore, the embedding layer necessitates frequent access to the off-chip memory system. As the embedding layer execution takes longer latency than the bottom MLP [11], the memory-bound nature of the embedding layer degrades the inference throughput.
@@ II-B NMP/PIM Designs for Recommendation System
Recent studies have integrated NMP/PIM units within DIMMs to accelerate the memory-intensive GnR operations. TensorDIMM [13] and RecNMP [12] have exploited rank-level NMP, whereas TRiM [22] leveraged bankgroup-level PIM. These units perform partial GnR on the data allocated within its node and transfer the result to the host. For embedding placement within memory, horizontal partitioning (HP) stores each embedding in a single memory node but suffers from load imbalance due to sparse, irregular access patterns. In contrast, TensorDIMM utilized vertical partitioning (VP), dividing an embedding across multiple ranks. VP is robust to the load imbalance but comes at the cost of energy inefficiency as concurrent row activations are incurred among multiple ranks. In pursuit of further speedup, the designs mentioned above leveraged the long-tail distribution of embeddings [11], [12] by placing a dynamic cache inside each NMP or introducing a load-balancing technique [12], [22]. SPACE enhanced throughput and power efficiency using HBM as a static DRAM cache for DIMM, storing frequently accessed embeddings and partial sums [11].
@@ III-A Weight-sharing Algorithms for Embedding Table Reduction
Weight-sharing algorithms, QR-trick [25] and TT-Rec [30], improve compression ratios and mitigate the accuracy loss of the hashing trick [28]. Weight-sharing algorithms transform an embedding table into several subtables consisting of subembeddings. Therefore, they require an additional reconstruction process before the GnR. This reconstruction, termed collect-and-reconst (CnR) in this paper, involves vector multiplication or GEMV operations.
QR-trick. The QR-trick uses double hashing with quotient and remainder hash functions to generate two subtables: the Q subtable and the R subtable, as shown in Figure 2 (a). The Q subtable maps consecutive original embeddings to the same subembedding based on identical quotient values, while the R subtable maps embeddings separated by the collision value based on identical remainder values. The R subtable size is negligible compared to the Q subtable, as it equals the hash collision value. During inference, the embedding layer reconstructs the original embedding from the subtables. Each hash function processes an input index to generate Q and R subtable indices for retrieving subembeddings. A mathematical operation is then performed on these subembeddings to reconstruct the embedding, where this paper focuses on element-wise vector multiplication, which generally outperforms across various models [25].
@@
TT-Rec. TT-Rec compresses embedding tables by decomposing them into tensor-train cores (TT-cores) through reshaping and decomposition, as shown in Figure 2 (b). The embedding table $T\in\mathbb{R}^{M\times N}$ is reshaped by factorizing $M$ and $N$ into $M=m1\times m2\times\cdots\times md$ and $N=n1\times n2\times\cdots\times nd$, resulting in $T\in\mathbb{R}^{(m1\times n1)\times(m2\times n2)\times\cdots\times(md\times nd)}$. Decomposition then converts the reshaped table into TT-cores, where kth TT-core shape is $C(k)=\mathbb{R}^{R_{k-1}\times(mk\times nk)\times R_{k}}$, $0<k<d$, where $R$ being the rank and $d$ the number of cores. The rank values are typically set to 16, 32, and 64 and total TT-cores to 3. Note that the TT-cores correspond to subtables in this work. The second subtable is three-dimensional; each entry consists of $R$ vectors. Reconstructing a full embedding is accomplished by fetching $n1$, $n2$, and $n3$ consecutive subembeddings from each subtable and then performing the GEMV followed by a dot product. TT-Rec results in a 327$\times$ reduction with the number of subtables and the rank value set to 3 and 32 [30].
@@ III-B Weight-Sharing Algorithms on Previous PIM Systems
The weight-sharing algorithm involves vector multiplication or GEMV operations in the embedding layer for CnR, which require vector pairs to reside in the same PIM unit for computation. However, it is impossible to co-locate all vector pairs within the same node because the subtables are distributed across multiple nodes (e.g., ranks, bank groups) for parallel processing. This requires CPU-PIM communication, as depicted in Figure 3 (a). The second CnR requires the memory controller to pass an operand in a different node. A similar incident occurs for the TT-Rec. Furthermore, as shown in Figure 3 (b), communication overhead takes more than 30% of the total execution time of the embedding layer for all embedding dimensions. Dynamic techniques to alleviate CPU-NMP communication, such as inter-node broadcast [27], face challenges with weight-sharing algorithms. The asynchronous nature of PIM operations for the embedding layer requires each PIM to fetch distinct data, resulting in idle cycles as PIM should wait for data over the shared bus. Although interconnect methods for DIMM-NMP [34] have been proposed to address point-to-point communication, modern PIM architectures still lack support for inter-PIM communication [10]. Even without communication, weight-sharing algorithms incur increased memory access as CnR requires access to all subtables. The abovementioned problems undermine the performance of prior PIM architectures. However, our roofline analysis [29] in Figure 4 shows that both algorithms are within the memory-bound region, indicating that PIM is still a promising solution for embedding layer with weight-sharing algorithms.
@@ III-C Access Patterns and Table Characteristics
Each subtable must be accessed concurrently in weight-sharing algorithms to reconstruct an original embedding. The total number of accesses is not distributed but remains the same for each subtable. In addition, the total subembeddings in a subtable are much smaller than the original embedding table. Thus, access to each subembedding becomes more concentrated than the original, intensifying the access rate. Another observation is that subembeddings are reused not only across GnRs of the same batch but also within a single GnR. In conventional recommendation systems, a user’s interaction with an item is typically reflected once in one sparse feature, preventing multiple accesses within a single GnR. However, weight-sharing algorithms map multiple embeddings to a single subembedding, necessitating a CnR before the GnR, thus enabling the reuse of subembeddings. We term this characteristic as intra-GnR locality. As shown in Figure 2 (b), access to consecutive embeddings cause multiple access to the single subembedding within the Q subtable. The intra-GnR locality is further strengthened for smaller subtable sizes, given that access to the subembeddings gets more concentrated. Previous designs are not optimized for such characteristics, losing an opportunity for extra speedup.
@@ IV ProactivePIM Architecture
We propose the ProactivePIM, a PIM architecture that supports weight-sharing algorithm. ProactivePIM employs an integrated approach for speedup by utilizing in-memory cache (detailed in Section IV-B), eliminating communication overhead through data mapping (detailed in Section IV-C), and minimizing additional overhead in the PIM system.
@@ IV-A ProactivePIM Architecture
A two-level PIM system is used inside HBM to leverage in-memory parallelism. In Figure 5 (b), the first-level processing unit (bd-PIM) is on the base die, while the second-level units (bg-PIM) are inside each bank group. Each bg-PIM performs CnR followed by GnR on the CnR outputs and passes the partial results to the bd-PIM for final reduction. Figure 5 (d) shows bg-PIM, where PIM instructions (PIM-Inst) delivered from the host are buffered and then decoded in the decoder to set control paths of the PIM unit. Memory-mapped register (MMReg) is placed within the decoder to indicate the prefetch address’s start and end. MAC unit performs CnR/GnR, while prefetched subembeddings are stored inside the cache to utilize intra-GnR locality. The HP method is employed for data placement, as its energy efficiency is better than VP’s. Figure 5 (e) shows bd-PIM, similar to the bg-PIM without the cache and MAC unit. The architecture can also support original DLRM or PageRank [7], leveraging cache to store feature vectors of frequently accessed vertices for higher throughput [31].
@@ IV-B Hot Subembedding Prefetch
To leverage intra-GnR locality, “hot” subembeddings with high access frequency are cached into bg-PIM before the CnR/GnR execution. During the CnR/GnR, hot subembedding accesses are directed to the cache. Partial sum transfer and prefetching are pipelined to minimize the prefetch overhead, by utilizing separate buses: transfer uses an external bus, while prefetching uses an internal bus, as shown in Figure 5 (d). When PIM-Inst for data transfer is issued from the decoder, bg-PIM starts to prefetch. To ensure correct timing for cached data access, the PIM extension needs to store physical addresses of cached data for reference. The corresponding area overhead is mitigated by prefetching an entire subtable. As described in Section IV-C, specific subtables are replicated in each bank group; the extension only checks the subtable ID for issuing cached data access as an entire subtable is prefetched.
Target Subtable to Prefetch. The key criteria for selecting a target subtable are (i) the degree of intra-GnR locality and (ii) the subtable size, both of which influence cache efficiency. In the QR-trick, the R subtable is much smaller than the Q subtable, and subembeddings exhibit high intra-GnR locality as described in Section III-A. Therefore, the R subtable is chosen. For TT-Rec, the intra-GnR locality is similar among all the subtables, but the second subtable is three-dimensional, resulting in a larger size. Furthermore, as described in Section IV-C, the subembedding dimension of the third table is smaller than the other tables, which degrades cache efficiency. Therefore, the first subtable is the target.
Prefetch Method. Figure 6 (a) shows the original execution timeline for CnR/GnR execution of three original tables. Once CnR/GnR completes for table A, the partial sums from each bg-PIM are transferred to the bd-PIM during the transfer phase. Figure 6 (b) illustrates the naive method in which all target subtables are read and offloaded before the embedding layer execution, which we name all-table prefetch. At the cost of the initial delay, the time consumed in PIM operation is largely reduced due to the cache hit. This effect is permanent before the inference terminates. However, this method requires a large SRAM cache to store all the target subtables. When there are 20 tables, 17% area overhead in DRAM die occurs, which worsens with more tables. VP could be utilized to reduce SRAM size fourfold by distributing subembedding dimensions across bank groups. However, for dimensions smaller than 256B, the data read per access falls below the DRAM burst granularity, reducing efficiency.
@@
As shown in equation 1, the transfer latency (tTransfer) is proportional to the total bank groups in a channel, read command delay between bank groups (tccd_s), batch size, and the embedding dimension (vlen). In contrast, the prefetch latency (tPrefetch) is proportional to total subembeddings within a target subtable (totsub), read command delay within bank groups (tccd_l), and subembedding dimension (subvlen). As a small batch size could cause severe load imbalance across PIM nodes, especially critical for a multi-stack HBM system, we choose a batch size 16. As described in Section VI-D, prefetch overhead is minimal, which indicates tPrefetch is similar to tTransfer. The host reads MMReg in the PIM to verify prefetch completion to enable next CnR/GnR execution. Target subtable data are mapped consecutively in a row to utilize the row buffer hit.